Question 1 Tries remaining: 2 Points out of 7.70 2 $ $10 Stax $8 Flag question $6 $4 $2 10 0) 30 40 50 Calculate the Deadweight Loss that this per unit tax would generate. (Do not include a S sign in your response. Round to the nearest two decimal places if necessary.) Answer: Check
Answer #1

$ $10 Stax $8 $4 DWL $2 10 20 30 40 50

Deadweight loss = 1/2 x base x height = 1/2 x (6 - 2) x (20 - 10) = 2 x 10 = $ 20